OpenStudy (anonymous):

Use trig substitution integral sqrt(x^2-1)/x^3 x=-2..-8 -2 is the lower, -8 is the upper, i started x = sec theta

OpenStudy (anonymous):

x^2 -1 = tan^2 theta, dtheta = sectheta tan theta

OpenStudy (anonymous):

i put it into my calculator and i get funny answers

OpenStudy (anonymous):

I get -.353, but the answer according to calculator is .353

OpenStudy (anonymous):

I have integral sqrt (tan^2 t ) / tan^3 t * sec t * tan t dt , where t = theta

OpenStudy (anonymous):

errr I have integral sqrt (tan^2 t ) / sec^3 t * sec t * tan t dt , where t = theta

OpenStudy (anonymous):

, so integral tan^2 / sec ^2 dt = integral sec^2 - 1 / (sec^2

OpenStudy (anonymous):

integral 1 - integral 1/sec^2

OpenStudy (anonymous):

integral 1 - integral cos^2

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sqrt{x^2-1}}{x^3} dx=\int\limits_{}^{}\frac{x}{x} \cdot \frac{\sqrt{x^2-1}}{x^3} dx\] \[=\int\limits_{}^{}\frac{x \cdot \sqrt{x^2-1}}{x^4} dx\] ----------------------------------------- Let \[u=\sqrt{x^2-1} => u^2=x^2-1 => 2 u du =2x dx => u du =x dx\] ----------------------------------------- \[=\int\limits_{}^{}\frac{u \cdot u}{(u^2+1)^2} du=\int\limits_{}^{}\frac{u^2}{(u^2+1)^2} du\] ------------------------------------------ \[\frac{u^2}{(u^2+1)^2}=\frac{Au+B}{u^2+1}+\frac{Cu+D}{(u^2+1)^2}\] \[=\frac{(Au+B)+(u^2+1)(Cu+D)}{(u^2+1)^2}\] \[=\frac{Au+B+Cu^3+Du^2+Cu+D}{(u^2+1)^2}\] => \[u^2=Au+B+Cu^3+Du^2+Cu+D=u^3(C)+u^2(D)+u(A+C)+(B+D)\] => \[C=0; D=1; A=0;B=-1\] ------------------------------------------ so we have \[\int\limits_{}^{}\frac{u^2}{(u^2+1)^2} du=\int\limits_{}^{} (\frac{-1}{u^2+1}+\frac{1}{(u^2+1)^2}) du\] \[=-\arctan(u)+\int\limits_{}^{}\frac{1}{(u^2+1)^2} du+C\]

OpenStudy (anonymous):

integral 1 - integral ( 1 + cos (2x)) / 2 t - t/2 - sin t *cos t / 2

OpenStudy (anonymous):

you should use trig substitution

myininaya (myininaya):

 \[\tan(\theta)=u => \sec^2(\theta) d \theta=du \] so we have \[\int\limits_{}^{}\frac{1}{(\tan^2(\theta)+1)^2} \sec^2(\theta) d \theta\]

OpenStudy (anonymous):

1/2 ( t - sin t * cos t ) , where t is theta 1/2 [ arcos (-1/8) - sin (arcos(-1/8)) cos (arcos(-1/8)] - arcos (-1/2) + sin (arcos -1/2) cos (arcos-1/2) )

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sec^2(\theta)}{\sec^4(\theta)} d \theta=\int\limits_{}^{}\cos^2(\theta) d \theta\] \[\frac{1}{2}\int\limits_{}^{}(1+\cos(2\theta)) d \theta =\frac{1}{2} (\theta+\frac{1}{2}\sin(2 \theta))+C\] \[=\frac{1}{2}(\theta+\frac{1}{2} \cdot 2 \sin(\theta) \cos(\theta))+C\] \[=\frac{1}{2} \theta+\frac{1}{2}\sin(\theta) \cos(\theta) +C\] \[=\frac{1}{2}\tan^{-1}(u)+\frac{1}{2} \cdot \frac{u}{\sqrt{u^2+1}} \cdot \frac{1}{\sqrt{u^2+1}}+C\] \[=\frac{1}{2}\tan^{-1}(u)+\frac{1}{2} \cdot \frac{u}{u^2+1}+C\] ------------------------------------------------------- so we have so far \[-\arctan(u)+\frac{1}{2}\arctan(u)+\frac{u}{2(u^2+1)}+C\] now we need to put terms of x

OpenStudy (anonymous):

can you check over my work , and im curious what you get as a final answer

myininaya (myininaya):

\[=-\arctan(\sqrt{x^2-1})+\frac{1}{2}\arctan(\sqrt{x^2-1})+\frac{\sqrt{x^2-1}}{2((x^2-1)+1)}+C\] and i guess we could combine are like terms \[=\frac{-1}{2}\tan^{-1}(\sqrt{x^2-1})+\frac{\sqrt{x^2-1}}{2x^2}+C\]

myininaya (myininaya):

\[\int\limits_{-8}^{-2}\frac{\sqrt{x^2-1}}{x^3} dx=[-\frac{1}{2}\tan^{-1}(\sqrt{x^2-1})+\frac{\sqrt{x^2-1}}{2x^2}]_{-8}^{-2}\] = \[[\frac{-1}{2}\tan^{-1}(\sqrt{4-1})+\frac{\sqrt{4-1}}{8}]\] - \[[-\frac{1}{2}\tan^{-1}(\sqrt{64-1})+\frac{\sqrt{64-1}}{128}\]

OpenStudy (anonymous):

my calculator gives me an answer of .3536 , but when i do it out i get -.3536 , im not sure what i did wrong

myininaya (myininaya):

= \[\frac{-1}{2}\tan^{-1}(\sqrt{3})+\frac{\sqrt{3}}{8}+\frac{1}{2}\tan^{-1}(\sqrt{63})-\frac{\sqrt{63}}{128}\]

myininaya (myininaya):

yes i get .3536

myininaya (myininaya):

did you do F(upper limit)-F(lower limit)

myininaya (myininaya):

oh wait the upper limit is -8?

myininaya (myininaya):

i thought it was the lower limit

myininaya (myininaya):

it would -.3536 then

OpenStudy (anonymous):

yes